OpenStudy (anonymous):

Let f(x) = \(\frac{x}{1+x}\) , x> -1. If \(g = f^{-1}\) , find the derivative of g using inverse function theorem.

OpenStudy (anonymous):

Inverse function theorem: Let f be differentiable and g be the inverse of f. If f'(c) \(\ne\) 0, then g is differentiable at d=f(c) and g'(d) = 1/ [f'(c)]. My attempt: \[f'(x) = \frac{1}{(1+x)^2}\]\[g= f^{-1}\]\[g’ = \frac{1}{f'(x)} =\frac{1}{\frac{1}{(1+x)^2}} = (1+x)^2\] Somehow, I’m quite sure that I didn’t use the theorem correctly. Can anyone show me the correct way or point out my mistakes Thanks in advance!

OpenStudy (anonymous):

It is a bit more complicated;

OpenStudy (anonymous):

1 Express , using only f(x) = y , the variable x as a function of its image y

OpenStudy (anonymous):

2 IN the formula u have obtained USE this expression x = some-formula(y) instead of x itself. Then simplify - and here is ur answer only expressed in y dummy variable. Change to x - and u have it

OpenStudy (anonymous):

Would you mind explaining step 1, particularly ''f(x) = y''?

OpenStudy (anonymous):

y = x/(1+x) ====> y + xy = x ===> x(1-y) = y

OpenStudy (anonymous):

x= y/(1-y)

OpenStudy (anonymous):

so plug this instead of x in YOU first answer

OpenStudy (anonymous):

So, did you mean express x in terms of y for the first step?

OpenStudy (anonymous):

y

OpenStudy (anonymous):

AND plug this instead of x in YOU first answer

OpenStudy (anonymous):

(1 + y/(1-y)) ^2

OpenStudy (anonymous):

How does it work? I'm sorry but I don't quite understand what you're doing.

OpenStudy (anonymous):

Have you computed the derivative of g ? - Yes you have Buuut this must be expressed in terms of the natural variable of g. This variable is y and NOT x

OpenStudy (anonymous):

Sooooo ... What should a poor studiosus do if he has an answer only in terms of x ?

OpenStudy (anonymous):

Eurica ! He should make an expression made entirely of y-dependence HWICH IS EQUIVALENT to x, and REPLACE each appearance of x with this expression ! (equal)

OpenStudy (anonymous):

That is what I recommended - and finally applied

OpenStudy (anonymous):

Actually, the answer given by the professor was g'(y) = 1/(1-y)^2. But he didn't explain how to get it though...

OpenStudy (anonymous):

Last small act - the letter used in the function is in some respect pure formality ... soooo..

OpenStudy (anonymous):

By the way, I don't understand how you find the derivative of g.... (I'm sorry again!)

OpenStudy (anonymous):

Wait let me first check YOUR original "solution"

OpenStudy (anonymous):

Correct

OpenStudy (anonymous):

Huh?! Correct?!

OpenStudy (anonymous):

Not in the sense that u should supply THAT as final expression

OpenStudy (anonymous):

Aaaand MY ANSWER IS IDENTICAL to your proff. answer

OpenStudy (anonymous):

Let us try to focus where're u loosing the thread - tell me

OpenStudy (anonymous):

You did the proper thing - but only half of needed transformation

OpenStudy (anonymous):

1. I really don't think I used that theorem correctly, but I don't know what's wrong with that. 2. I don't understand why it is in terms of y in the answer given. Yes, I know they are identical but I'm really confused.

OpenStudy (anonymous):

The "puenta" , essence of the answer is to use th variable y, not the variable x

OpenStudy (anonymous):

The answer is given in y-s and NOT in x-s , is because g is formally allowed to act (compute) only upon y-s.

OpenStudy (anonymous):

So - you must express each x in ur answer (correct, but using the wrong variable) by the expression of y(x)

OpenStudy (anonymous):

Sorry by x(y)

OpenStudy (anonymous):

This is what I did.

OpenStudy (anonymous):

To give u a comparable meta-fore:

OpenStudy (anonymous):

Suppose h(x) = x^5, then obviously h'(x) = 5*x^4

OpenStudy (anonymous):

Now what if someone would supply u with the following instructions:

OpenStudy (anonymous):

"LET z = x^2, and I TELL you truly, that h'(x) = 5*z^2

OpenStudy (anonymous):

What would you notice ? And what would you correct by substitution ?

OpenStudy (anonymous):

Firstly U would notice that his notation is a bit inconsistnet : h'9OF X) = Z-SQUARED

OpenStudy (anonymous):

He has two different variables , ond z is NOT the right variable

OpenStudy (anonymous):

You, would "correct" this by substituting z = x^2

OpenStudy (anonymous):

Similar substitution was done HERE

OpenStudy (anonymous):

I supposed \(g = f^{-1}\) is \(g(x) = f^{-1}(x)\)... Is this the mistake I made?

OpenStudy (anonymous):

Formally speaking this is NOt a mistake. Your ONLY mistake is not COMPLETING. The COMPLETION means using in the final expression the variable of g and not of f

OpenStudy (anonymous):

The problem is why g the function in y not in x. I don't understand this part.

OpenStudy (anonymous):

Why is it NOT formally a mistake you ask? Because any variable letter can be replaced by another - they are dummy vars. g(z). g(t), g(y) are all the same function !

OpenStudy (anonymous):